OpenStudy (adunb8):

how do you solve for x^3 - x^2 +3x+2 ?

OpenStudy (anonymous):

solve what?

OpenStudy (adunb8):

solve for x

OpenStudy (anonymous):

you do not have an equation \(x\) can be any number

OpenStudy (adunb8):

x^3-x^2+3x+2 = 0 and solve for what x needs to be

OpenStudy (anonymous):

you cant factor it. there is a way to solve for x, which is an extremely complicated process that i would never have been able to do, but the answer can be found here: http://www.wolframalpha.com/input/?i=x%5E3-x%5E2%2B3x%2B2+%3D+0

OpenStudy (adunb8):

hm..... i remember someone doing taking parenthesis both sides like (x^3-x^2) -(3x+2) = 0 or something anyone ring a bell?

OpenStudy (anonymous):

that wont work. for that to work, when you factor the 2 individual equations, they must have something in common after you factor them

OpenStudy (anonymous):

@ adunb8 what subject are u in?

OpenStudy (adunb8):

this is linear algebra

OpenStudy (adunb8):

/ differential eqn

OpenStudy (adunb8):

um i dont think so..

OpenStudy (adunb8):

sorry i got the sign wrong. its y'''-y''-3y'+2y=0

OpenStudy (adunb8):

it doesn't say i need to find derivative but yeah if i do it then its 3x^2-2x-3+0

OpenStudy (adunb8):

you got the sign wrong its -3 for the last number but yea that is y'

OpenStudy (anonymous):

are you trying to solve it?

OpenStudy (anonymous):

so this is the correct prob? x^3 - x^2 -3x+2 ?

OpenStudy (adunb8):

yea

OpenStudy (anonymous):

ok do you know synthetic division?

OpenStudy (adunb8):

i totally forgot =(

OpenStudy (anonymous):

doing your synthetic division try +1,-1,2,-2,3,-3

OpenStudy (adunb8):

what i dont understand what +1,-1,2,-2,3,-3 are....

OpenStudy (anonymous):

ok if you have x^3 - x^2 -3x+2 =0 try 2/ 1 -1 -3 2 2 2 -2 ----------------- 1 1 -1 0 ok that was perpect there for x=2 that is y=0= x^3 - x^2 -3x+2 =(x-2)(x^2 +x -1) here one root of x=2,, now try to find the other two roots

OpenStudy (anonymous):

Refer to a solution using the Mathematica program.